# CAOS: Conformal Aggregation of One-Shot Predictors

### v1

@@ 6 Experiments
We evaluate CAOS as a method for producing calibrated prediction sets for one-shot prediction in comparison to several Split Conformal One-shot Predictor (SCOS) baselines, namely (i) exemplar-based facial landmarking with pretrained vision models, and (ii) Real World Few-Shot Annotated Tasks (RAFT) with LLMs.
@@ 6.1 Exemplar-Based Facial Landmarking
While pretrained vision models provide a strong foundation for landmark detection, fully supervised fine-tuning is often impractical in this setting. Clinical annotations are expensive and time-consuming, and while hundreds of labeled examples may be feasible, large-scale annotation is rarely realistic. As a result, approaches that rely on extensive fine-tuning are typically out of reach.
Instead, each annotated image can be viewed as an exemplar that induces a one-shot predictor, for example by transferring landmarks via patch similarity in a pretrained embedding space. Aggregating such predictors provides a natural way to leverage limited labeled data without additional training. However, exemplar relevance varies substantially across target images, making uncertainty-aware adaptive aggregation essential for reliable deployment.
We evaluate CAOS in this regime using publicly available face data, enabling controlled evaluation while supporting reproducibility and dissemination of results.
Facial Landmarking Dataset.  We evaluate exemplar-based facial landmarking using aligned face images from the CelebA dataset. For reproducibility, the exact hash of the dataset version we used along with other relevant implementation details is provided in Appendix D.1.
@@
Ground truth landmark locations are obtained via the MediaPipe Face Landmarker with default settings333num_faces=1, min_face_detection_confidence=0.5, min_face_detection_confidence=0.5, which places 478 landmarks in each image. This auto-annotation scheme provides a scalable way to generate labeled data for 478 landmarking tasks of varying difficulty. While we ackgnowledge that MediaPipe Face Landmarker could make mistakes, it nevertheless produces a meaningful benchmark for the evaluation of one-shot landmarking via patch similarity.
@@
Approaches for Conformal One-Shot Landmarking.  Given a dataset of $|\mathcal{D}_{n}|$ labeled examples, we can instantiate $|\mathcal{D}_{n}|$ one-shot predictors. Rather than point predictions, we want to obtain conformal prediction sets at a target miscoverage level $\alpha$, which are guaranteed to contain the true label with finite-sample marginal probability at least $1-\alpha$.
Our guiding research question is how to leverage the available data to combine the one-shot predictors (e.g. by selecting the best ones for a new test example) while also using the available data to calibrate the conformal prediction sets.
We evaluate CAOS in comparison to several split conformal baselines in terms of empirical coverage and prediction set size (Eqns. 32 and 33 in Appendix D.2). Empirical coverage should meet or exceed the target level $1-\alpha$, while smaller prediction sets, corresponding to lower predictive uncertainty, are preferred. For a fair comparison, all methods use the same one-shot prediction tasks, backbone models, and data splits, differing only in how prediction sets are constructed and calibrated.
  • SCOS Avg.: This baseline follows Alg. 2 to calibrate $|\mathcal{D}_{\textrm{ref}}|$ one-shot predictors on calibration data $\mathcal{D}_{\mathrm{cal}}$. Performance is reported on average over one-shot predictors.
@@
  • CAOSours: Our approach uses $\mathcal{D}_{n}$ both as the pool of reference examples and as calibration data according to Alg. 1. To avoid issues with hyperparameter selection we set the number of smallest scores to aggregate $k=3$ throughout.
@@
Figure 1: Prediction set size per landmark for exemplar-based facial landmarking. Each point corresponds to a landmark. CAOS yields smaller prediction sets than SCOS Avg. and SCOS Best, approaching oracle efficiency.
@@
ResultsTable 1 summarizes results in terms of empirical coverage and average set size at different target miscoverage rates $\alpha\in[0.05,0.1,0.2]$. We report mean and standard error around the mean averaged over the held-out test data and the 487 landmarking tasks. CAOS consistently meets or exceeds the target coverage rate, while significantly improving over the SCOS prediction sets in terms of size.
The scatter plot in Figure 1 further breaks down these results at $\alpha=0.05$ by individual landmarks which vary significantly by difficulty; on some landmarks even the oracle that selects the smallest prediction set that contains the true label achieves only prediction set sizes that exceed 20 patches on average (averaged over held-out set). The x-axis is the prediction set size achieved by the SCOS oracle, while the values on the y-axis give the corresponding set size achieved by the other methods, SCOS Avg. (gray circles) SCOS Best (blue triangles) and CAOS (red diamonds).
Figure 2 visualizes the one-shot prediction results for 478 facial landmarks in 3 test images. We color each ground truth landmark by the set size achieved SCOS Avg. (top) and CAOS (bottom) (the smaller the better, values below 10 are marked in green). We can see qualitatively that CAOS achieves significantly smaller prediction set sizes.
@@ 6.2 Real-World Few-Shot Annotated Tasks with LLMs
We evaluate CAOS on the RAFT benchmark, a collection of real-world few-shot text classification tasks designed to reflect practical annotation workflows. Each task provides a small labeled dataset and exhibits substantial heterogeneity in label space size, class balance, and semantic complexity, making RAFT a challenging testbed for uncertainty quantification in low-data regimes.
We treat each labeled example as inducing a one-shot predictor via ICL with a fixed prompt template. Given a reference example and a test input, a large language model produces a candidate label score without task-specific fine-tuning. This setup naturally yields a collection of one-shot predictors per task, whose quality varies across inputs and labels.
Our goal is to construct conformal prediction sets that adaptively aggregate these one-shot predictors while using the limited labeled data efficiently for calibration. We compare CAOS against split conformal one-shot baselines under identical backbones, prompts, and data splits, differing only in how prediction sets are calibrated and aggregated.
@@
One-shot Text Classification with LLMs.  We evaluate CAOS on one-shot text classification tasks from the RAFT benchmark, following the setup described in Sec. 3, Example 2, using a fixed prompt template (see Sec. D.3) and a frozen pretrained language model, specifically a 7B-parameter model from the Llama2 family.
Each RAFT task provides 50 labeled examples. We construct a dataset $\mathcal{D}_{n}$ of size 40 and a held-out test set $\mathcal{D}_{\textrm{test}}$ of size 10, which is used exclusively to evaluate empirical coverage and prediction set size. For split conformal baselines, which require disjoint reference and calibration data, $\mathcal{D}_{n}$ is further partitioned into equally sized reference and calibration sets $\mathcal{D}_{\text{ref}}$ and $\mathcal{D}_{\text{calib}}$. All methods use identical data splits, prompts, and backbone models, differing only in how prediction sets are constructed and calibrated.
For each candidate label $y$ in the task-specific label space, we form a prompt by concatenating the reference example, the test input, and the candidate output. We compute the negative log-likelihood (NLL) of $y$ under the model, and use the length-normalized NLL as the nonconformity score to avoid penalizing longer labels. While we use Llama2-7B in our experiments, this procedure applies to any language model that provides access to token-level logits, including API-based models, provided the label space is known.
Results.Figure 3 summarizes the performance of CAOS and SCOS on one-shot text classification tasks from the RAFT benchmark using Llama2-7B at $\alpha=0.1$. Across tasks, CAOS consistently produces smaller prediction sets than SCOS, achieving lower average set size in 8 out of 9 tasks. When both methods satisfy the target coverage level (4/9 tasks), CAOS maintains comparable coverage while yielding tighter prediction sets. In more challenging regimes, where SCOS fails to meet the target coverage, CAOS remains valid in 3 additional tasks. Only in 2 tasks do both methods fail to achieve the desired coverage. [^4] Each RAFT task has a small evaluation set of size 10, which makes empirical coverage estimates inherently noisy. At a target coverage level of $1-\alpha=0.9$, even a perfectly calibrated method may observe only 8 out of 10 test examples covered in a given split. Accordingly, occasional deviations from the target coverage should be expected in this low-sample regime. These results indicate that CAOS improves efficiency without sacrificing reliability in the one-shot setting.
@@ 7 Conclusion
Despite breaking classical exchangeability assumptions at the score level, we proved that CAOS achieves exact finite-sample marginal coverage. Our analysis relies on a novel monotonicity-based argument that reduces the CAOS procedure to full conformal prediction, thereby avoiding the slack terms that typically arise in cross-conformal or data-adaptive settings. This result shows that algorithmic departures from exchangeability need not preclude sharp coverage guarantees when carefully structured aggregation rules are used.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Notation Summary
This appendix summarizes the notation used throughout the paper.
Data and indices.
  • $\mathcal{X}$: input space.
  • $\mathcal{Y}$: output (label) space.
  • $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$: labeled reference dataset.
  • $(X_{n+1},Y_{n+1})$: test example with ground truth label $Y_{n+1}$.
  • $i,j$: indices of reference examples.
  • $y\in\mathcal{Y}$: candidate output label.
One-shot predictors and nonconformity scores.
  • $s_{\pi_{i}}(x,y)=-\pi_{i}(y\mid x)$: one-shot nonconformity score for target $(x,y)$ using predictor $\pi_{i}$.
Aggregation and Quantile operators.
  • $\operatorname{Quantile}\!\left(\mathcal{S}\,;\,\tau\right)$: the $\lceil\tau(|\mathcal{S}|+1)\rceil$-th order statistic of the multiset $\mathcal{S}$, for $\tau\in(0,1)$.
            This definition accounts for ties and ensures finite-sample validity.
Split conformal, CAOS, and full CAOS nonconformity scores.
  • Test-time scores
    • $s_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n}\bigr)=\frac{1}{k}\Sigma^{k}_{\min}(\mathcal{A}_{\mathcal{D}_{n}}\!\left(X_{n+1},y\right))$: CAOS test-time aggregated nonconformity score.
    • $\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n+1}^{y}\bigr)=\frac{1}{k}\bigl(\Sigma^{k+1}_{\min}(\mathcal{A}_{\mathcal{D}_{n+1}^{y}}\!\left(X_{i},Y_{i}\right))-s_{\pi_{i}}(X_{i},Y_{i})\bigr)$: full CAOS test-time nonconformity score.
  • Calibration scores
    • $s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr)=s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr)$: leave-one-out CAOS calibration score for example $i$.
    • $\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)=\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)$: full CAOS calibration score for example $i$.
  • Relationships between nonconformity scores
    • $\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n+1}^{y}\bigr)=s_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n}\bigr)$: test-time score equivalence between CAOS and full CAOS under Assumption 4.1.
Quantiles and prediction sets.
  • $\hat{q}_{\mathrm{caos}}$: CAOS calibration threshold.
  • $\hat{q}_{\mathrm{full}}^{y}$: full CAOS calibration threshold.
  • $\hat{C}_{\mathrm{caos}}(X_{n+1})$: CAOS prediction set.
  • $\hat{C}_{\mathrm{full}}(X_{n+1})$: full conformal CAOS prediction set.
Appendix B Pseudocode
@@
Appendix C Detailed Proofs for CAOS Theory
C.1 Proof of Test-Time Score Equivalence
We prove the identity
@@ C.1 Proof of Test-Time Score Equivalence
which is used in the set-inclusion argument in Sec. 5.
Recall that the CAOS test-time nonconformity score is defined as
$s_{\mathrm{caos}}\bigl((X_{n+1},y);\mathcal{D}_{n}\bigr)=\frac{1}{k}\,\Sigma^{k}_{\min}\!\bigl(\{s_{\pi_{i}}(X_{n+1},y):(X_{i},Y_{i})\in\mathcal{D}_{n}\}\bigr).$ (28)
For full CAOS, the augmented dataset is $\mathcal{D}_{n+1}^{y}=\mathcal{D}_{n}\cup\{(X_{n+1},y)\}$, and the corresponding test time score is
$\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y);\mathcal{D}_{n+1}^{y}\bigr)=\frac{1}{k}\Bigl(\Sigma^{k+1}_{\min}\!\bigl(\mathcal{A}_{\mathcal{D}_{n+1}^{y}}(X_{n+1},y)\bigr)-s_{\pi_{n+1}}(X_{n+1},y)\Bigr),$ (29)
where $\mathcal{A}_{\mathcal{D}_{n+1}^{y}}(X_{n+1},y)$ denotes the multiset of one-shot nonconformity scores obtained by using each element of $\mathcal{D}_{n+1}^{y}$ as a reference for the target $(X_{n+1},y)$.
By construction,
$\mathcal{A}_{\mathcal{D}_{n+1}^{y}}(X_{n+1},y)=\{s_{\pi_{i}}(X_{n+1},y):(X_{i},Y_{i})\in\mathcal{D}_{n}\}\;\cup\;\{s_{\pi_{n+1}}(X_{n+1},y)\}.$ (30)
Under Assump. 4.1, the self-score $s_{\pi_{n+1}}(X_{n+1},y)$ is minimal among all one-shot scores for the target $(X_{n+1},y)$. Consequently, it is always included among the $(k+1)$ smallest elements of $\mathcal{A}_{\mathcal{D}_{n+1}^{y}}(X_{n+1},y)$.
Subtracting this self-score therefore removes exactly one of the selected $(k+1)$ terms, leaving the sum of the $k$ smallest scores induced by $\mathcal{D}_{n}$. Dividing by $k$ yields
$\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y);\mathcal{D}_{n+1}^{y}\bigr)=\frac{1}{k}\,\Sigma^{k}_{\min}\!\bigl(\{s_{\pi_{i}}(X_{n+1},y):(X_{i},Y_{i})\in\mathcal{D}_{n}\}\bigr),$ (31)
which is precisely $s_{\mathrm{caos}}\bigl((X_{n+1},y);\mathcal{D}_{n}\bigr)$. ∎
Appendix D Additional Implementation Details
D.1 Facial Landmarking Data

### v2

@@ 6 Experiments
We evaluate CAOS as a method for producing calibrated prediction sets for one-shot prediction in comparison to several Split Conformal One-shot Predictor (SCOS) baselines on (i) exemplar-based facial landmarking with pretrained vision models, and (ii) Real World Few-Shot Annotated Tasks (RAFT) with LLMs (Alex et al., 2021).
@@
No marginal coverage guarantee due to model selection on the calibration data; included only as a reference.
@@ 6.1 Exemplar-Based Facial Landmarking
Clinical annotations are expensive and time-consuming and large-scale annotation is rarely realistic. With a few labeled examples, each annotated image can be used to induce a one-shot predictor. Since relevance varies substantially across predictors, uncertainty-aware adaptive aggregation is essential for reliable deployment. We evaluate CAOS using publicly available face data to enable controlled and reproducible evaluation.
Facial Landmarking Dataset.  We evaluate exemplar-based facial landmarking using aligned face images from the CelebA dataset. For reproducibility, the exact hash of the dataset version we used along with other relevant implementation details is provided in Appendix C.1.
@@
Ground truth landmark locations are obtained via the MediaPipe Face Landmarker with default settings111num_faces=1, min_face_det_confidence=0.5, which places 478 landmarks in each image. This auto-annotation scheme provides a scalable way to generate labeled data for 478 landmarking tasks of varying difficulty. While we acknowledge that MediaPipe Face Landmarker could make mistakes, it nevertheless produces a meaningful benchmark for the evaluation of one-shot landmarking via patch similarity.
@@
Approaches for Conformal One-Shot Landmarking.  We evaluate CAOS in comparison to several split conformal baselines in terms of empirical coverage and prediction set size (Eqns. 28 and 29 in Appendix C.2). Empirical coverage should meet or exceed the target level $1-\alpha$, while smaller prediction sets, corresponding to lower predictive uncertainty, are preferred. For a fair comparison, all methods use the same one-shot prediction tasks, backbone models, and data splits, differing only in how prediction sets are constructed and calibrated.
  • SCOS: This baseline follows Alg. 2 to calibrate $|\mathcal{D}_{\textrm{ref}}|$ one-shot predictors on calibration data $\mathcal{D}_{\mathrm{cal}}$. Performance is reported on average over one-shot predictors.
@@
  • CAOSours: CAOS uses $\mathcal{D}_{n}$ both as the pool of reference examples and as calibration data according to Alg. 1.
Choice of $k$:  In principle, the aggregation parameter $k$ could be selected via cross-validation. However, in the one-shot setting considered here, labeled data are extremely scarce, and cross-validation would require withholding examples that could otherwise be used for reference or calibration. We therefore fix $k=3$ for all experiments and empirically verify that CAOS performance is robust to the choice of $k$ across a wide range of values (App. D).
@@
Figure 1: Prediction set size per landmark for exemplar-based facial landmarking. Each point corresponds to a landmark. CAOS yields smaller prediction sets than SCOS and SCOS Best, approaching oracle efficiency.
@@
ResultsTable 1 summarizes results in terms of empirical coverage and average set size at different target miscoverage rates $\alpha\in[0.05,0.1,0.2]$. We report mean and standard error around the mean averaged over the held-out test data and the 478 landmarking tasks. CAOS consistently meets or exceeds the target coverage rate, while significantly improving over the SCOS prediction sets in terms of size.
The scatter plot in Figure 1 further breaks down these results at $\alpha=0.05$ by individual landmarks which vary significantly by difficulty; on some landmarks even the oracle that selects the smallest prediction set that contains the true label achieves only prediction set sizes that exceed 20 patches on average (averaged over held-out set). The x-axis is the prediction set size achieved by the SCOS oracle, while the values on the y-axis give the corresponding set size achieved by the other methods, SCOS (gray circles) SCOS Best (blue triangles) and CAOS (red diamonds).
Figure 2 visualizes the one-shot prediction results for 478 facial landmarks in 3 test images. We color each ground truth landmark by the set size achieved by SCOS (top) and CAOS (bottom) (the smaller the better, values below 10 are marked in green). We can see qualitatively that CAOS achieves significantly smaller prediction set sizes.
To isolate the source of these gains, App. E presents a diagnostic ablation showing that aggregation alone yields limited improvements when data are split, and that reusing labeled examples for both calibration and reference is the main driver for CAOS ’s efficiency. Thm. 4.1 is essential to establish that such data reuse is valid.
@@ 6.2 Real-World Few-Shot Annotated Tasks with LLMs
We evaluate CAOS on the RAFT benchmark (Alex et al., 2021), a collection of real-world few-shot text classification tasks designed to reflect practical annotation workflows. Each task provides a small labeled dataset and exhibits substantial heterogeneity in label space size, class balance, and semantic complexity, making RAFT a challenging testbed for uncertainty quantification in low-data regimes.
@@
One-shot Text Classification with LLMs.  Each RAFT task provides 50 labeled examples. Each example induces a one-shot predictor via ICL with a fixed prompt template (see Sec. C.3), as described in Sec. 3, Ex. 2. For each candidate label $y$ in the task-specific label space, we concatenate the label with the prompt and compute the negative log-likelihood (NLL) of $y$ under the model, and use the length-normalized NLL as the nonconformity score to avoid penalizing longer labels. While we use Llama2-7B in our experiments, this procedure applies to any language model that provides access to token-level logits, including API-based models, provided the label space is known.
We construct a dataset $\mathcal{D}_{n}$ of size 40 and a held-out test set $\mathcal{D}_{\textrm{test}}$ of size 10, which is used exclusively to evaluate empirical coverage and prediction set size. For split conformal baselines, which require disjoint reference and calibration data, $\mathcal{D}_{n}$ is further partitioned into equally sized reference and calibration sets $\mathcal{D}_{\text{ref}}$ and $\mathcal{D}_{\text{calib}}$. All methods use identical data splits, prompts, and backbone model, differing only in how prediction sets are constructed and calibrated.
Results.Figure 3 summarizes the performance of CAOS and SCOS on one-shot text classification tasks from the RAFT benchmark using Llama2-7B at $\alpha=0.1$. Across tasks, CAOS consistently produces smaller prediction sets than SCOS, achieving lower average set size in 8 out of 9 tasks. When both methods satisfy the target coverage level (4/9 tasks), CAOS maintains comparable coverage while yielding tighter prediction sets. In more challenging regimes, where SCOS fails to meet the target coverage, CAOS retains coverage in 3 additional tasks. Only in 2 tasks do both methods fail to achieve the desired coverage. [^2] Each RAFT task has an evaluation set of size 10, which makes empirical coverage estimates inherently noisy. At a target coverage level of $0.9$, the probability of observing $8\leq 10$ covered is about $26\%$ under a Binomial(1010,$0.9$) model. Accordingly, occasional misscoverage should be expected in this low-sample regime. These results indicate that CAOS reduces predictive uncertainty and improves efficiency when labeled data is scarce.
@@ 7 Conclusion
Despite breaking classical exchangeability assumptions at the score level, we proved that CAOS achieves finite-sample marginal coverage. Our analysis relies on a novel monotonicity-based argument that reduces the CAOS procedure to full conformal prediction, thereby avoiding the slack terms that typically arise in cross-conformal or data-adaptive settings. This result shows that algorithmic departures from exchangeability need not preclude sharp coverage guarantees when carefully structured aggregation rules are used.
@@
This work develops a methodological contribution to uncertainty quantification in machine learning, with a focus on conformal prediction in one-shot and low-data regimes. The proposed method is intended to improve the statistical efficiency of prediction sets when labeled data are scarce, which is a common setting in scientific and medical applications.
However, as with all uncertainty quantification methods, the established guarantees apply only under the stated assumptions, and misuse or misinterpretation of prediction sets could lead to unwarranted confidence. The method does not introduce new data collection, deployment, or automation mechanisms, and we do not anticipate specific negative societal impacts beyond those generally associated with the application domains in which conformal prediction is used.
Appendix A Pseudocode
@@
Appendix B Detailed Proofs for CAOS Theory
B.1 Proof of Symmetry of the full CAOS score
Proof of Lemma 5.1.
The operator $\Sigma^{k}_{\min}$ depends only on the multiset of its inputs and is therefore symmetric. Moreover, the subtracted term in Eqn. 18 corresponds to the one-shot score induced by the target example itself and is unaffected by permutations of the remaining elements of $\mathcal{D}$. It follows that $\tilde{s}_{\mathrm{caos}}((x,y);\mathcal{D})=\tilde{s}_{\mathrm{caos}}((x,y);\mathcal{D}_{\sigma}).$
B.2 Proof of the Monotonicity of the CAOS score
Proof of Lemma 5.2.
$\Sigma^{k}_{\min}(\mathcal{A}^{\prime})=\sum_{j=1}^{k}a^{\prime}_{(j)}\;\leq\;\sum_{j=1}^{k}a_{(j)}=\Sigma^{k}_{\min}(\mathcal{A}),$
yielding $s_{\mathrm{caos}}((x,y);\mathcal{D}^{\prime})\leq s_{\mathrm{caos}}((x,y);\mathcal{D}),$ as claimed. ∎
B.3 Score comparison between CAOS and full CAOS
Fix a candidate label $y\in\mathcal{Y}$ and the augmented dataset $\mathcal{D}_{n+1}^{y}=\mathcal{D}_{n}\cup\{(X_{n+1},y)\}$. Our goal is to show that the test scores are equivalent
@@ B.3 Score comparison between CAOS and full CAOS
while for the calibration scores indexed by $i=1,\ldots,n$,
Proof of Lemma 5.3.
Both statement follow from the definitions of the scores and the definition of multiset substraction provided in Eqn. 30, App. F.
For the test inputs we have
$\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y);\mathcal{D}_{n+1}^{y}\bigr)$ $=\Sigma^{k}_{\min}\bigl(\mathcal{A}_{\mathcal{D}_{n+1}^{y}}\!\left(X_{n+1},y\right)-\{s_{\pi_{n+1}^{y}}(X_{n+1},y)\}\bigr)$
$=\Sigma^{k}_{\min}\bigl(\mathcal{A}_{\mathcal{D}_{n}}\!\left(X_{n+1},y\right)\bigr)$
$=s_{\mathrm{caos}}\bigl((X_{n+1},y);\mathcal{D}_{n}\bigr).$ (26)
Similarly, for the calibration scores:
$\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)$ $=\Sigma^{k}_{\min}\bigl(\mathcal{A}_{\mathcal{D}_{n+1}^{y}}\!\left(X_{i},Y_{i}\right)-\{s_{\pi_{i}}(X_{i},Y_{i})\}\bigr)$
$=\Sigma^{k}_{\min}\bigl(\mathcal{A}_{\mathcal{D}_{n}^{-i}\cup\{(X_{n+1},y)\}}(X_{i},Y_{i})\bigr)$
$=s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\cup\{(X_{n+1},y)\}\bigr)$
$\leq s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr).$ (27)
The last inequality follows from Lemma 5.2 (Monotonicity Lemma). ∎
Appendix C Additional Implementation Details
C.1 Facial Landmarking Data